# The GECKOS Survey: Identifying kinematic sub-structures in edge-on galaxies

### v1

@@ 2.3 Data Reduction
The dedicated python package pymusepipe v2.23.4  [^3] https://github.com/emsellem/pymusepipe was utilised to calibrate the science exposures and create mosaicked data cubes. pymusepipe is a nearly fully automated pipeline, described in detail in Emsellem et al. (2022). It serves as a data organiser and wrapper for the sequential execution of the ESO Recipe Execution Tool (esorex; ESO CPL Development Team, 2015) recipes, and also provides additional utilities for improved alignment, flux and sky calibration, and mosaicking. pymusepipe is built around the MUSE Data Reduction Pipeline (DRP) that was developed to remove instrumental signatures (Weilbacher et al., 2020) [^4] https://www.eso.org/sci/software/pipelines/muse/. A full description of the GECKOS data reduction workflow will be presented in van de Sande et al. (in prep.).
In this paper, we use all galaxies whose observations were completed by December 2023, for either the entire galaxy or the central pointing. This subset of galaxies forms the core of the GECKOS first internal data release (iDR1) and includes fully reduced and mosaicked data cubes for the galaxies ESO 484-036, NGC 5775, ESO 079-003, UGC 00903, ESO 120-016, NGC 3279, NGC 0360, PGC 044931, IC 1711, and NGC 3957, and the central pointing only of NGC 0522 and IC 5244. This sub-sample contains galaxies from all three bins of star formation rate. Three-colour $g^{\prime}ri$ images reconstructed from the MUSE cubes of the 12 galaxies used in this paper are shown in Figure 1. We note here that while the MUSE nominal wavelength range of $\sim$4800–9300 Å  covers the SDSS $r$ and $i$ bands well, the $g$ band is truncated, and so we refer to it as $g^{\prime}$. The reconstructed $g^{\prime}$ band image follows the SDSS $g$-band filter response curve for $\sim$4800–5500 Å.
@@ 2.4 Data Analysis
There exist many pipelines for the analysis of IFS data, several of which wrap existing well-established spectral fitting routines, for example, the data analysis pipelines of MaNGA (Westfall et al., 2019) and PHANGS (Emsellem et al., 2022), Pipe3D (Sánchez et al., 2016) and the GIST pipeline (Bittner et al., 2019). The future of extragalactic, spatially-resolved spectroscopy will involve pushing current instrumentation to its limits, requiring highly customisable analysis software packages. The GECKOS Survey alone necessitates the development of techniques to deal with low S/N, and highly sky-dominated regions to extract science from the galactic outer disc regions. Additionally, large co-added mosaics of several pointings of nearby galaxies are becoming a common technique to resolve the small-scale physics driving galaxy evolution. Such mosaics are large, and computationally expensive to run through existing, non-optimised software. Here, we present nGIST [^5] https://github.com/geckos-survey/ngist, an upgraded version of the GIST  pipeline, with added features, improved usability, and retained functionality and core principles. Described more fully in Appendix A, nGIST is publicly available, documented, and maintained via a GitHub repository.
@@
Following the recommendations outlined in the pPXF code documentation, we determined the optimized bias value by performing a large ensemble of Monte Carlo simulations, testing how well the LOSVD parameters are recovered as a function of velocity dispersion. The ideal bias is defined as one that reduces the scatter in the velocity dispersion, $h_{3}$ and $h_{4}$, without creating a systematic offset in the velocity and velocity dispersion when the dispersion is larger than the spectrum’s velocity scale (41 km s-1). We used a MILES SSP with an age of 5.01 Gyr, [M/H] = 0.0, and a wavelength range of 4800–7000Å.
@@ 3 Results
Figure 2: $r$-band DECaLS image (top) scaled with an inverse hyperbolic sine ($\operatorname{arcsinh}$) scaling with surface brightness contours overlaid in white (units of $r$-band magnitude arcsec-2), and unsharp masked image created with a Gaussian smoothing kernel of width $\sigma=7-20$ pixels (bottom). Galaxy names marked with a $\star$ are considered to host a boxy-peanut central structure in this work.
@@ 3.2 Visual morphologies from the literature
Table 1: Summary of morphological classifications from unsharp masking of DECaLS imaging (this work) and previous literature. A ‘–’ denotes that a galaxy was not included in a particular sample.
@@
    Digitised Sky Survey imaging; pixel scale 1.7′′ pixel-1.
@@
3.3 Stellar kinematic maps
@@
In Figures 314, we present the derived stellar line-of-sight luminosity-weighted mean velocity ($V_{\star}$), velocity dispersion ($\sigma_{\star}$), and high-order moments skew ($h_{3}$) and kurtosis ($h_{4}$) for the GECKOS galaxies, derived over the wavelength range 4800–8900 Å. We also tested just the Calcium II triplet (CaT) range (8450–8735 Å, similar to McDermid, 2002), but found that while the redder wavelength range resulted in more spatially coherent structure in the high-order moment maps, this short wavelength range often failed to produce a cohesive $h_{3}$ and $h_{4}$ signal in the outer, lower dispersion regions of some of the galactic discs. For a further discussion of wavelength-dependent kinematic structures, see Appendix B. In Figures 314 each galaxy is rotated such that the dust lane (and photometric major axis) lie horizontal and we note that for some galaxies, the MUSE mosaic covers only one side of the disc. The GECKOS pointing strategy was such that the entirety of the bulge was included in the MUSE pointings, however. At first glance, all galaxies appear to host regularly rotating discs. However, upon more careful examination, the diversity in structure in the kinematic maps is considerable.
3.4 1D radial profiles
@@
For a more straightforward comparison to previous literature, Figures 15 and 16 depict the 1D radial profiles extracted from the corresponding 2D maps. Depending on the dust lane structure, these profiles were extracted from the midplane regions of each galaxy in slits of width 0.2–0.8 kpc. To aid in the recognition of kinematic structure, we have flipped the direction of the profiles of ESO 120-016, NGC 0360, NGC 0522, ESO 484-036, and IC 5244 such that the receding side of the resultant rotation curve is on the right for all galaxies. Below, we describe the trends seen in the 2D kinematic maps and 1D radial profiles, and in Section 4 we interpret these results in terms of galactic sub-structures.
3.4.1 Velocity, $V_{\star}$, maps
The top left panels of Figures 314 display $V_{\star}$ maps, corrected for systemic velocity using a circular aperture of 1 kpc centred on the flux centre of the galaxy. In all cases, we see evidence of well-behaved, regularly-rotating discs, with the notable exception of UGC 00903, which appears to comprise two counter-rotating stellar discs (see van de Sande et al. in prep. for further insight). In some cases, the dust lane is evident as a lighter region just beneath the midplane. Some galaxies show ‘twist’-like features along their kinematic minor axes (e.g. ESO 079-003, NGC 0522, and NGC 0360). We expect that these features are due to the fact that the depth at which we are peering into the galaxy varies as a function of position within the galaxy. Near the midplane, where the dust is thickest, our sight-line is more impeded than in the outer, more dust-free regions.
@@ 3.4.1 Velocity, $V_{\star}$, maps
For IC 1711, NGC 3957, and PGC 044931, the most striking feature is a small, darker-coloured, central region that is rotating faster than the surrounding disc, likely a kinematically-decoupled discy component. For comparison to previous literature, in Figure 15a, we plot the galaxy rotation curves, extracted from slits of width 0.2–0.8 kpc (depending on dust geometry) centered about the midplane. All galaxies host rapidly rising rotation curves, though the slope of the curve is much steeper in the central regions of PGC 044931, NGC 3957, IC 1711, and IC 5244. The characteristic ‘double-humped’ rotation curve described by Bureau & Athanassoula (2005) is seen in these four galaxies.
@@
Figure 3: Stellar luminosity-weighted line-of-sight velocity, $V_{\star}$, (top left), velocity dispersion, $\sigma_{\star}$, (top right, with $r$-band surface brightness contours overlaid in white), $h_{3}$ (bottom left) and $h_{4}$ (bottom right) maps derived with nGIST and binned to S/N=100 for the galaxy NGC 3279, a galaxy with no obvious visual bulge structure. Galaxies are shown such that the dust lane (and photometric major axis) are oriented horizontally. Holes in the maps are regions where bright stars have been masked. Flux contours are overlaid on the $\sigma_{\star}$ maps to highlight the dust lane effects, but not for the $V_{\star}$, $h_{3}$ and $h_{4}$ maps, to aid in central structure recognition.
@@
Figure 8: As for Figure 3, but for NGC 0360 - no obvious visual bulge structure.
@@
Figure 11: As for Figure 3, but for UGC 00903 - no obvious visual bulge structure.
Figure 12: As for Figure 3, but for NGC 5775 - no obvious visual bulge structure.
@@
Figure 15: 1D line profiles, extracted from the midplane regions of each galaxy but avoiding the dust lane, from slits of width 0.2–0.8 kpc, depending on dust position. Each coloured point represents a Voronoi bin, and the white line is a running median. To enable easier comparisons, in all cases, the galaxy is rotated such that the receding side is to the right. Typical median errors are $V_{\star,err}\sim 2$ km s-1 and $\sigma_{\star,err}\sim 5$ km s-1
@@
3.4.2 Stellar velocity dispersion, $\sigma_{\star}$, maps
@@
In addition to the dust lanes, we also observe central regions of high-velocity dispersion, transitioning to lower dispersion at greater radii in most of the galaxies, except NGC 3279 and UGC 00903. Central $\sigma_{\star}$ maxima range from 70–130 km s-1, with an average of $\sigma_{\star}\sim 90$ km s-1. The increase in $\sigma_{\star}$ from the outskirts to the central regions, $\Delta\sigma_{\star}$, is on average $\sim 50$ km s-1. In NGC 3957 and PGC 044931, this central high $\sigma_{\star}$ region is ‘croissant-shaped’, with a lower dispersion region embedded within the high $\sigma_{\star}$ region at the centre of the galaxy. We expect that these are not the result of dust obscuration, as neither of these galaxies is perfectly edge-on and the effect of the dust is visible at lower latitudes. Additionally, the low $\sigma_{\star}$ regions are detached from the main dust lane; co-located with the kinematically-decoupled disc indicator seen in the $V_{\star}$ maps. For IC 5244 and IC 1711, the central region with $\sigma_{\star}>100$ km s-1 is extended and nearly circular, with dispersions exceeding 110 km s-1.
@@
3.4.3 $h_{3}$ maps
The bottom left panels of Figures 314 show the high-order Gauss-Hermite moment $h_{3}$ (or skew). The colour map is such that red indicates a positive skew value (where excess receding velocities are present in the LOSVD beyond that of a pure Gaussian), and blue a negative (where excess approaching velocities are present in the LOSVD). While there is more stochasticity between neighbouring bins than the $V_{\star}$ and $\sigma_{\star}$ maps, some spatially-coherent features are readily visible.
We observe small, oval-shaped regions of highly positive and negative $h_{3}$ at the very centre of the galaxy and only along the midplane for IC 5244, IC 1711, NGC 3957, and PGC 044931. Compared to their $V_{\star}$ maps, these regions are anti-correlated with velocity. We expect these features to be the signature of a dynamically cold disc, and discuss them further in Section 4.2.
At larger scales, for some galaxies we also observe coherent positive (negative) regions of $h_{3}$, corresponding to positive (negative) values of $V_{\star}$, extending 2–5 kpc radially, and 1–2 kpc off the midplane of the disc. These structures in $h_{3}$ correlate with the sign of $V_{\star}$ in the velocity maps. We observe the strongest $h_{3}$-$V_{\star}$ correlation regions in IC 5244, and IC 1711, but also see this feature to varying degrees in all other BP bulge galaxies, though the signal is faint in ESO 120-016 and NGC 0522. In the outskirts of all galaxies in this sample, we note coherent regions of $h_{3}$-$V_{\star}$ anti-correlation, extending to the edges of the MUSE field of view. For galaxies with the most $h_{3}$ structure, moving from the central regions to the galactic outskirts, we observe a $h_{3}$-$V_{\star}$ anti-correlation, correlation, and then anti-correlation again.

### v2

@@ 2.3 Data reduction
The dedicated python package pymusepipe v2.23.4  [^4] https://github.com/emsellem/pymusepipe was utilised to calibrate the science exposures and create mosaicked data cubes. pymusepipe is a nearly fully automated pipeline, described in detail in Emsellem et al. (2022). It serves as a data organiser and wrapper for the sequential execution of the ESO Recipe Execution Tool (esorex; ESO CPL Development Team, 2015) recipes, and also provides additional utilities for improved alignment, flux and sky calibration, and mosaicking. pymusepipe is built around the MUSE Data Reduction Pipeline (DRP) that was developed to remove instrumental signatures (Weilbacher et al., 2020) [^5] https://www.eso.org/sci/software/pipelines/muse/. A full description of the GECKOS data reduction workflow will be presented in van de Sande et al. (in prep.), and we note that a similar setup was used in Watts et al. (2024).
Three-colour $g^{\prime}ri$ images reconstructed from the MUSE cubes of the 12 galaxies used in this paper are shown in Figure 1. We note here that while the MUSE nominal wavelength range of $\sim$4800–9300 Å  covers the SDSS $r$ and $i$ bands well, the $g$ band is truncated, and so we refer to it as $g^{\prime}$. The reconstructed $g^{\prime}$ band image follows the SDSS $g$-band filter response curve for $\sim$4800–5500 Å.
@@ 2.4 Data analysis
There exist many pipelines for the analysis of IFS data, several of which wrap existing well-established spectral fitting routines, for example, the data analysis pipelines of MaNGA (Westfall et al., 2019) and PHANGS (Emsellem et al., 2022), Pipe3D (Sánchez et al., 2016) and the GIST pipeline (Bittner et al., 2019). The future of extragalactic, spatially-resolved spectroscopy will involve pushing current instrumentation to its limits, requiring highly customisable analysis software packages. The GECKOS Survey alone necessitates the development of techniques to deal with low S/N, and highly sky-dominated regions to extract science from the galactic outer disc regions. Additionally, large co-added mosaics of several pointings of nearby galaxies are becoming a common technique to resolve the small-scale physics driving galaxy evolution. Such mosaics are large, and computationally expensive to run through existing, non-optimised software. Here, we present nGIST, an upgraded version of the GIST  pipeline, with added features, improved usability, and retained functionality and core principles. Described more fully in Appendix A, nGIST is publicly available, documented, and maintained via a GitHub repository.
@@
Following the recommendations outlined in the pPXF code documentation, we determined the optimized bias value by performing a large ensemble of Monte Carlo simulations, testing how well the LOSVD parameters are recovered as a function of velocity dispersion. The ideal bias is defined as one that reduces the scatter in the velocity dispersion, $h_{3}$ and $h_{4}$, without creating a systematic offset in the velocity and velocity dispersion when the dispersion is larger than the spectrum’s velocity scale (41 km s-1). In the tests, we used a MILES SSP with an age of 5.01 Gyr, [M/H] = 0.0, and a wavelength range of 4800–7000Å.
@@ 3 Results
Figure 2: $r$-band DECaLS image (top) scaled with an inverse hyperbolic sine ($\operatorname{arcsinh}$) scaling with surface brightness contours overlaid in white (units of $r$-band magnitude arcsec-2), and unsharp masked image created with a Gaussian smoothing kernel of width $\sigma=7-20$ pixels (bottom). Galaxy names marked with a $\star$ are considered to host a boxy-peanut central structure in this work. The orientation of galaxies here, and throughout, is the same as in Fig 1.
@@ 3.2 Visual morphologies from the literature
Table 1: Summary of galaxy morphological classifications from the literature.
@@
    Digitised Sky Survey imaging; pixel scale 1.7′′ pixel-1.  [^7] Classifications from this work come from the unsharp masking of DECaLS imaging. A ‘–’ denotes that a galaxy was not included in a particular sample.
@@
3.3 Stellar kinematic maps and radial profiles
@@
In Figures 314, we present the derived stellar line-of-sight luminosity-weighted mean velocity ($V_{\star}$), velocity dispersion ($\sigma_{\star}$), and high-order moments skew ($h_{3}$) and kurtosis ($h_{4}$) for the GECKOS galaxies, derived over the wavelength range 4800–8900 Å. We also tested just the Calcium II triplet (CaT) range (8450–8735 Å, similar to McDermid, 2002), but found that while the redder wavelength range resulted in more spatially coherent structure in the high-order moment maps, this short wavelength range often failed to produce a cohesive $h_{3}$ and $h_{4}$ signal in the outer, lower dispersion regions of some of the galactic discs. For a further discussion of wavelength-dependent kinematic structures, see Appendix B. In Figures 314 each galaxy is rotated such that the photometric major axis lies horizontal and the dust lane (if offset), is always to the lower side of the image. We note that for some galaxies, the MUSE mosaic covers only one side of the disc. The GECKOS pointing strategy was such that the entirety of the bulge was included in the MUSE pointings, however. At first glance, all galaxies appear to host regularly rotating discs. However, upon more careful examination, the diversity in structure in the kinematic maps is considerable.
For a more straightforward comparison to previous literature, Figures 15 and 16 depict the 1D radial profiles extracted from the corresponding 2D maps. These profiles were extracted from slits of width 1 kpc between $-0.3<\rm{z}<0.7$, which we found to be wide enough to encompass central kinematic structure of interest, whilst minimising the contribution of dust lane regions, keeping in mind that the dust lanes are always offset to the lower side of the images for which we place the slits on. Furthermore, rather than presenting $h_{3}$ directly against $V_{\star}$, we present all 1D kinematic quantities as a function of galactocentric radius. To facilitate interpretation, for 1D profiles we compare the radial gradients of $V_{\star}$ and $h_{3}$, where if both gradients are positive or negative, the quantities are correlated, but if one is positive and the other is negative, the two quantities are anti-correlated. For the 2D maps, we discuss the sign of the $V_{\star}$ and $h_{3}$ regions (i.e. a positive value of $V_{\star}$ and $h_{3}$ have a matching sign).
@@
To aid in the recognition of kinematic structure, we have flipped the direction of the profiles of ESO 120-016, NGC 0360, NGC 0522, ESO 484-036, IC 1711, and IC 5244 such that the receding side of the resultant rotation curve is on the right for all galaxies. Below, we describe the trends seen in the 2D kinematic maps and 1D radial profiles, and in Section 4 we interpret these results in terms of galactic sub-structures.
3.3.1 Velocity, $V_{\star}$, maps and profiles
The top left panels of Figures 314 display $V_{\star}$ maps, corrected for systemic velocity using a box of width 2 kpc extending perpendicular to the galaxy kinematic major axis, centred on the flux centre of the galaxy, and extending to only the side of the galaxy that is less dust affected (i.e. positive values of z in Figures 314). We mask regions with E(B-V)¿0.4, and recalculate the median velocity within this region before subtracting it from the maps. In all cases, we see evidence of well-behaved, regularly-rotating discs, with the notable exception of UGC 00903, which appears to comprise two counter-rotating stellar discs (see van de Sande et al. in prep. for further insight). In some cases, the dust lane is evident as a region lighter in colour, indicating velocity magnitude lower than its surroundings, located just below the z=0 line. For many galaxies the $V_{\star}=0$ isovelocity line is not straight; rather, often warps, wedges, or other structures are visible both along the z=0 line, and at higher latitudes. For several galaxies (e.g. PGC 044931, NGC 0522, NGC 0360, and ESO 079-003), the kinematic minor axis displays a wedge-like structure within ±$\pm 0.5$kpc of the z=0 line. The co-spatial nature of young stars and dust make light-weighted stellar kinematic maps difficult to interpret in highly obscured regions, though we speculate on the commonly observed features in Section 4.1.2.
@@ 3.3.1 Velocity, $V_{\star}$, maps and profiles
For IC 1711, NGC 3957, and PGC 044931, the most striking feature is a small, darker-coloured region within the central kiloparsec that is rotating faster than the surrounding disc, likely a kinematically-decoupled discy component. For comparison to previous literature, in Figure 15a, we plot the galaxy rotation curves, extracted from slits of width 1 kpc. All galaxies host rapidly rising rotation curves, though the slope of the curve is much steeper in the central regions of PGC 044931, NGC 3957, IC 1711, and IC 5244. The characteristic ‘double-humped’ rotation curve (i.e. a rotation curve that first rises rapidly to reach a local maximum, then drops slightly to create a local minimum, and then rises again slowly up to its flat section) described by Bureau & Athanassoula (2005) is seen in these four galaxies.
@@
Figure 3: Stellar luminosity-weighted line-of-sight velocity, $V_{\star}$, (top left), velocity dispersion, $\sigma_{\star}$, (top right, with $r$-band surface brightness contours overlaid in white), $h_{3}$ (bottom left) and $h_{4}$ (bottom right) maps derived with nGIST and binned to S/N=100 for the galaxy NGC 3279, a galaxy with no obvious visual bulge structure. Galaxies are shown such that the dust lane (and photometric major axis) are oriented horizontally, and such that the dust lane (if offset) is to the lower side of the image. Holes in the maps are regions where bright stars have been masked. Flux contours are overlaid on the $\sigma_{\star}$ maps to highlight the dust lane effects, but not for the $V_{\star}$, $h_{3}$ and $h_{4}$ maps, to aid in central structure recognition.
@@
Figure 8: As for Figure 3, but for NGC 0360 - no obvious visual BP bulge structure.
@@
Figure 11: As for Figure 3, but for UGC 00903 - no obvious visual BP bulge structure.
Figure 12: As for Figure 3, but for NGC 5775 - no obvious visual BP bulge structure.
@@
Figure 15: 1D line profiles, extracted from slits of width the $0.70<\rm{z}<$ 0.7 kpc. Each coloured point represents a Voronoi bin, and the white line is a running median. To enable easier comparisons, in all cases, the galaxy is rotated such that the receding side is to the right. Typical median errors are $V_{\star,err}\sim 2$ km s-1 and $\sigma_{\star,err}\sim 5$ km s-1
@@
3.3.2 Stellar velocity dispersion, $\sigma_{\star}$, maps and profiles
@@
In addition to the dust lanes, we also observe central regions of high-velocity dispersion, transitioning to lower dispersion at greater radii in most of the galaxies, except NGC 3279 and UGC 00903. Central $\sigma_{\star}$ maxima range from 70–130 km s-1, with an average of $\sigma_{\star}\sim 90$ km s-1. The increase in $\sigma_{\star}$ from the outskirts to the central regions, $\Delta\sigma_{\star}$, is on average $\sim 50$ km s-1. In NGC 3957 and PGC 044931, this central high $\sigma_{\star}$ region is ‘croissant-shaped’, with a lower dispersion region embedded within the high $\sigma_{\star}$ region at the centre of the galaxy. We expect that these are not the result of dust obscuration, as neither of these galaxies is perfectly edge-on and the effect of the dust is visible below the z=0 line. Additionally, the low $\sigma_{\star}$ regions are offset from the projection of the main dust lane in the 2D maps; co-located with the kinematically-decoupled disc indicator seen in the $V_{\star}$ maps. For IC 5244 and IC 1711, the central region with $\sigma_{\star}>100$ km s-1 is extended and nearly circular, with dispersions exceeding 110 km s-1.
@@
3.3.3 Skew, $h_{3}$, maps and profiles
The bottom left panels of Figures 314 show the high-order Gauss-Hermite moment $h_{3}$ (or skew). The colour map is such that red indicates a positive skew value (where excess low velocities are present in the LOSVD beyond that of a pure Gaussian), and blue a negative (where excess high velocities are present in the LOSVD). While there is more stochasticity between neighbouring bins than the $V_{\star}$ and $\sigma_{\star}$ maps, some spatially-coherent features are readily visible.
We observe small, oval-shaped regions of highly positive and negative $h_{3}$ at the very centre of the galaxy and only along the z=0 line for IC 5244, IC 1711, NGC 3957, and PGC 044931. Compared to their $V_{\star}$ maps, these regions are opposite in sign. Additionally, this feature is also seen in the 1D radial profiles of Figure 16a. A strong, negative gradient in $h_{3}$ as a function of galactic radius is seen in the $h_{3}$ profiles of PGC 044931, NGC 3957, IC 1711, IC 5244, and UGC 00903. For these galaxies, the extrema of the $h_{3}$ coincide radially with the $V_{\star}$ humps pointed out in Section 3.3.1. The two $h_{3}$ extrema present close to the galaxy centre imply large gradient in the central $h_{3}$ radial 1D profile between them, and since the signs of these extrema are opposite to those of $V_{\star}$, the corresponding gradient of the $V_{\star}$ radial profile is of opposite sign and therefore $h_{3}$ and $V_{\star}$ anti-correlate within the radius of the $h_{3}$ extrema. We expect these features to be the signature of a dynamically cold disc, and discuss them further in Section 4.2.
At larger scales, we see varying h3 behaviour along the z=0 line compared to off the plane for a given R in several galaxies: NGC 3957, NGC 0522, PGC 044931, IC 5244, and perhaps best illustrated in ESO 079-003. We observe $\sim 0.5$ kpc-wide strips where $h_{3}$ and $V_{\star}$ are of opposite sign along the kinematic major axis of these galaxies, discussed in Section 4.1.3. These strips often extend to the radial extent of the MUSE field of view. We note that the correlations seen in 1D profiles often extend only to small scales (e.g. $\sim$1 kpc for IC 1711), whereas when sign-matching, additional coherent regions of sign matching and mismatching are also obvious at larger scales. For IC 1711, both above and below the z=0 line, these opposing signs in $V_{\star}$ and $h_{3}$ switch to matching signs, and we observe coherent positive (negative) regions of $h_{3}$, corresponding to positive (negative) values of $V_{\star}$, extending 0–3 kpc radially, and 0.5–2 kpc off the z=0 line of the disc. We see the largest regions of matching $h_{3}$ and $V_{\star}$ sign in IC 5244, and IC 1711, but also see this feature to varying degrees in all other BP bulge galaxies, though the signal is faint in ESO 120-016 and NGC 0522. In the radial outskirts of all galaxies in this sample, we note coherent regions of $h_{3}$ and $V_{\star}$ opposing signs, extending to the edges of the MUSE field of view. For galaxies with the most coherent $h_{3}$ structure, moving from the central regions to the galactic outskirts, we observe the signs of $h_{3}$ and $V_{\star}$ to oppose, match, and then oppose again.